Solve -3(2w – 1) = 5w – 19. Please show all of your work. CAN SOMEONE SHOW ME THE STEPS TO DO THIS SO I DONT HAVE TO ASK...PLZ SOMEONE PLZ...!!!
Answer by checkley79(3341)

-3(2w – 1) = 5w – 19 expand the -3(2w-1) terms.
-6w+3=5w-19 combine the w terms & the integers.
-6w-5w=-19-3 add like terms.
-11w=-22 divide both sides by -11.
w=-22/-11
w=2 ans.
Proof:
-3(2*2-1)=5*2-19
-3(4-1)=10-19
-3*3=-9
-9=-9